OpenStudy (anonymous):

A beam of electrons is caused to move in a circular path of radius 3.00 m at a velocity of 2.00 x 10^7 m/s. The electron mass is 9.11 x 10^-13 kg. What is the centripetal force on one electron?

OpenStudy (anonymous):

a=v^2/R=>F=m*v^2/R F=121 metres